# Epistemic Uncertainty in Conformal Scores: A Unified Approach

### v1

@@ 4.2 Regression baselines
For regression-based conformal scores, we use a neural network optimized with a penalized Mean Squared Loss. Detailed descriptions of the architectures and hyperparameters used can be found in Appendix B.3.
We compare EPICSCORE to the following baselines:
@@ 4.3 Results
The mean average coverage is close to the nominal 90% for all methods (Table 5 and 8 of Appendix B.2), which is expected since all methods are conformal.
In the quantile regression setting, Table 1 highlights EPICSCORE’s strong performance across all datasets, with the MDN-MC Dropout variant excelling in 12 out of 13 cases. Notably, EPICSCORE outperforms all competitors on 7 datasets, effectively balancing coverage and interval precision. Additionally, Tables 6 and 7 (Appendix B.2) show that the MDN-MC Dropout variant produces narrower predictive intervals while maintaining good conditional coverage.
For regression, Table 2 shows EPICSCORE’s strong performance across most datasets. The MDN-MC Dropout variant excels in 12 out of 13 cases, while the BART version ranks among the top in 8 datasets. Additionally, Tables 9 and 10 (Appendix B.2) confirm its effectiveness in interval length, as well as approximate conditional coverage. These results highlight EPICSCORE’s flexibility and robustness in capturing epistemic and aleatoric uncertainty across different conformal scores.
@@ 5 Final Remarks
Looking ahead, we will extend EPICSCORE to settings with distribution shift, where capturing epistemic uncertainty is crucial due to data sparsity in some regions. By refining our approach for domain adaptation, we aim to maintain reliable predictive regions even when test distributions differ from training data.
Code to implement EPICSCORE and reproduce the experiments is available at https://github.com/Monoxido45/EPICSCORE.
@@ Acknowledgements.
L.M.C.C is grateful for the fellowship provided by São Paulo Research Foundation (FAPESP), grant 2022/08579-7. V. S. S. is grateful for the financial support from FAPESP (grant 2023/05587-1). R. I. is grateful for the financial support of FAPESP (grants 2019/11321-9 and 2023/07068-1) and CNPq (grants 422705/2021-7 and 305065/2023-8).
@@ A.1 Regression
In this section, we detail the example presented by Figure 1. We simulate a scenario with two distinct dense regions exhibiting low aleatoric and epistemic uncertainty, separated by an intermediate, sparser region with high aleatoric and epistemic uncertainty. Given a sample size $n$, we first generate $\left\lfloor{0.425\cdot n}\right\rfloor$ samples for each $X\sim U(0,1.5)$ and $X\sim U(8,10)$, ensuring that $85\%$ of the data comes from the two outer regions, reflecting low epistemic uncertainty. The corresponding response variable follows $Y\sim N(2\sin{X},0.1)$, which also reflects low aleatoric uncertainty. For the remaining $\left\lfloor{0.425\cdot n}\right\rfloor$ samples, we draw $X$ from a transformed Beta distribution, $X\sim(\text{Beta}(8,8)\cdot(8-1.5))+1.5$, concentrating points in the intermediate region. Here, the response variable follows $Y\sim N(2\sin{X},2.1)$, introducing high aleatoric uncertainty. Epistemic uncertainty is particularly elevated at the boundaries of this region.
In this setting, we use a K-nearest neighbors (KNN) algorithm with $k=10$ as the regression base model $g(\mathbf{x})$. We illustrate the difference between EPICSCORE and established conformal prediction baselines for regression intervals, including Regression Split [Lei and Wasserman, 2014], Weighted Regression Split [Lei et al., 2018], and the Mondrian Conformal Regression [Boström and Johansson, 2020]. For EPICSCORE, we use the BART-based version, with $m=100$ trees, default prior options for all parameters, and a heteroscedastic gamma distribution as the probability model for the conformal score (detailed in C.4), which is appropriate given that the regression conformal score is non-negative and often asymmetric.
@@ A.2 Classification
For the image classification example (introduced in Figure 2), we used the publicly available CIFAR-100 dataset [Krizhevsky et al., 2009], which consists of $60,000$ color images of size $32\times 32$ spanning $100$ classes, with each class containing $600$ images. The dataset was split into training, calibration, and test sets, allocating $10\%$ for testing and $45\%$ each for training and calibration. We utilized a ResNet-34 model [He et al., 2016] to extract 512-dimensional feature representations and trained a Random Forest classifier on the training set with default parameters except for the number of trees, which we fixed at $300$. The classifier reports an accuracy of $47\%$ in the test set. The Adaptive Prediction Set and EPICSCORE were then applied to the calibration set, using the MDN MC-dropout variant of EPICSCORE, which model’s architecture is better detailed in C.2.
@@
These results show that EPICSCORE-MDN achieves an SSC much closer to the nominal level of $0.8$ than APS, indicating more consistent coverage across different set sizes. In contrast, APS has a lower average SSC, reflecting greater deviations from the target coverage and potentially less reliable uncertainty quantification.
@@
To differentiate inliers from outliers for further analysis and comparison of each method’s prediction sets, we first apply t-SNE [Van der Maaten and Hinton, 2008] to reduce the dimensionality of the feature space in the test set. We then use the Local Outlier Factor (LOF) method [Breunig et al., 2000] for outlier detection, leveraging a KNN-based density estimation to identify anomalies. The LOF score not only highlights outliers but also helps characterize typical (inlier) observations, offering a structured approach to assessing epistemic uncertainty across different regions of the data distribution. We fit LOF on the first two t-SNE dimensions, assuming a contamination rate of $10\%$ of the sample is contaminated (i.e., treating $10\%$ of the sample as outliers), and rank the top 150 outliers and inliers for analysis.
@@
Appendix B Additional Results and Details for Real Data Experiments
@@
B.1 Evaluation Metrics
Let $R(\cdot)$ denote a generic prediction interval. Given a test set $(\mathbf{X}_{1},Y_{1}),(\mathbf{X}_{2},Y_{2}),\ldots,(\mathbf{X}_{m},Y_{m})$, we evaluate performance using the following metrics:
@@ B.1 Evaluation Metrics
    $\text{AMC}=\frac{1}{m}\sum_{i=1}^{m}\mathbb{I}\left(Y_{i}\in R(\mathbf{X}_{i})\right),$
@@
    where $\min\hat{R}(\mathbf{X})$ and $\max\hat{R}(\mathbf{X})$ represent the lower and upper bounds of the prediction interval, respectively, and $\alpha$ is the miscalibration level. The Interval Score Loss balances two key objectives: maintaining narrow prediction intervals while penalizing those that fail to cover $Y_{i}$, with larger penalties for greater deviations. By averaging these scores across all instances, AISL provides a measure that prioritizes the shortest interval while ensuring sufficient coverage.
@@
    where $\mathbf{C}=(C_{1},\ldots,C_{m})$ represents a binary vector, with $C_{i}=\mathbb{I}(Y_{i}\in R(\mathbf{X}_{i}))$ indicating whether the prediction interval $R(\mathbf{X}_{i})$ covers $Y_{i}$, and $\mathbf{W}=(W_{1},\ldots,W_{m})$, where $W_{i}=\max\hat{R}(\mathbf{X}_{i})-\min\hat{R}(\mathbf{X}_{i})$. A strong correlation between coverage and interval width suggests a potential violation of conditional coverage, which requires their independence [Feldman et al., 2021]. However, $\rho=0$ does not guarantee conditional coverage, as non-adaptive methods like regression split can achieve zero correlation by maintaining constant-width intervals [Rossellini et al., 2024]. Thus, while this metric provides a useful proxy for assessing conditional coverage, it is not a definitive measure.
B.2 Additional results
The dataset details are provided in Table 4. The evaluation results for quantile regression are summarized in Tables 5, 6, and 7, while the corresponding results for standard regression are presented in Tables 8, 9, and 10.
@@ B.2 Additional results
B.3 Base model hyperparameters

### v2

@@ 4.2 Regression baselines
For regression-based conformal scores, we use a neural network optimized with a penalized Mean Squared Loss. Detailed descriptions of the architectures and hyperparameters used can be found in Appendix C.3. We compare EPICSCORE to the following baselines:
@@ 4.3 Results
The mean average coverage is close to the nominal 90% for all methods (Table 6 and 10 of Appendix C.2), which is expected since all methods are conformal.
In the quantile regression setting, Table 1 highlights EPICSCORE’s strong performance across all datasets, with the MDN-MC Dropout variant ranking among the top in 12 out of 13 cases. Notably, EPICSCORE outperforms all competing methods on 7 datasets, effectively balancing coverage and interval sharpness. Table 3 demonstrates EPICSCORE’s superior adaptability to data-sparse regions, with the MDN-MC Dropout variant achieving the highest outlier-to-inlier interval length ratio in 10 out of 13 cases. Table 9 further confirms good coverage for outliers in those same sparse regions. Additionally, Tables 7 and 8 (Appendix C.2) show that this variant yields narrower intervals while maintaining strong conditional coverage.
For the regression setting, Table 2 confirms EPICSCORE’s strong performance across most datasets, with the MDN-MC Dropout variant excelling in 12 out of 13 cases and the BART version ranking among the top in 8. Tables 13 and 14 report the outlier-to-inlier interval length ratio and outlier coverage, respectively—both showing that the MDN-MC Dropout variant performs consistently well in data-sparse regions. Additionally, Tables 11 and 12 (Appendix C.2) confirm its ability to generate concise intervals while preserving approximate conditional coverage.
@@ 5 Final Remarks
Looking ahead, we plan to extend EPICSCORE to settings involving distribution shift, where accounting for epistemic uncertainty is important due to increased data sparsity in certain regions. In addition, we intend to improve the efficiency of our calibration procedure by eliminating the need to split the calibration set. To this end, we will explore adaptations based on Jackknife+ and CV+ methods [Barber et al., 2021], as well as multiple-split conformal prediction techniques [Lei et al., 2018], which offer ways to aggregate information with statistical guarantees. Code to implement EPICSCORE and reproduce the experiments is available at https://github.com/Monoxido45/EPICSCORE.
@@ Acknowledgements.
This study was financed in part by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brasil (CAPES) - Finance Code 001. L.M.C.C is grateful for the fellowship provided by São Paulo Research Foundation (FAPESP), grant 2022/08579-7. V. S. S. is grateful for the financial support of FAPESP (grant 2023/05587-1). R. I. is grateful for the financial support of FAPESP (grants 2019/11321-9 and 2023/07068-1) and CNPq (grants 422705/2021-7 and 305065/2023-8). The authors are also grateful to Rodrigo F. L. Lassance for his suggestions and insightful discussions.
@@ A.1 Regression
In this section, we detail the example presented by Figure 1. We simulate a scenario with two distinct dense regions exhibiting low aleatoric and epistemic uncertainty, separated by an intermediate, sparser region with high aleatoric and epistemic uncertainty. Given a sample size $n$, we first generate $\left\lfloor{0.425\cdot n}\right\rfloor$ samples for each $X\sim U(0,1.5)$ and $X\sim U(8,10)$, ensuring that $85\%$ of the data comes from the two outer regions, reflecting low epistemic uncertainty. The corresponding response variable follows $Y\sim N(2\sin{X},0.1)$, which also reflects low aleatoric uncertainty. For the remaining $\left\lfloor{0.15\cdot n}\right\rfloor$ samples, we draw $X$ from a transformed Beta distribution, $X\sim(\text{Beta}(8,8)\cdot(8-1.5))+1.5$, concentrating points in the intermediate region. Here, the response variable follows $Y\sim N(2\sin{X},2.1)$, introducing high aleatoric uncertainty. Epistemic uncertainty is particularly elevated at the boundaries of this region.
In this setting, we use a K-nearest neighbors (KNN) algorithm with $k=10$ as the regression base model $g(\mathbf{x})$. We illustrate the difference between EPICSCORE and established conformal prediction baselines for regression intervals, including Regression Split [Lei and Wasserman, 2014], Weighted Regression Split [Lei et al., 2018], and the Mondrian Conformal Regression [Boström and Johansson, 2020]. For EPICSCORE, we use the BART-based version, with $m=100$ trees, default prior options for all parameters, and a heteroscedastic gamma distribution as the probability model for the conformal score (detailed in D.4), which is appropriate given that the regression conformal score is non-negative and often asymmetric.
@@ A.2 Classification
For the image classification example (introduced in Figure 2), we used the publicly available CIFAR-100 dataset [Krizhevsky et al., 2009], which consists of $60,000$ color images of size $32\times 32$ spanning $100$ classes, with each class containing $600$ images. The dataset was split into training, calibration, and test sets, allocating $10\%$ for testing and $45\%$ each for training and calibration. We utilized a ResNet-34 model [He et al., 2016] to extract 512-dimensional feature representations and trained a Random Forest classifier on the training set with default parameters except for the number of trees, which we fixed at $300$. The classifier reports an accuracy of $47\%$ in the test set. The Adaptive Prediction Set and EPICSCORE were then applied to the calibration set, using the MDN MC-dropout variant of EPICSCORE, which model’s architecture is better detailed in D.2.
@@
These results show that EPICSCORE-MDN achieves an SSC much closer to the nominal level of $0.8$ than APS, indicating more consistent coverage across different set sizes. In contrast, APS has a lower average SSC, reflecting greater deviations from target coverage and potentially less reliable uncertainty quantification.
@@
To differentiate inliers from outliers for further analysis and comparison of each method’s prediction sets, we first apply t-SNE [Van der Maaten and Hinton, 2008] to reduce the dimensionality of the feature space in the test set. We then use the Local Outlier Factor (LOF) method [Breunig et al., 2000] for outlier detection, leveraging a KNN-based density estimation to identify anomalies. The LOF score not only highlights outliers but also helps characterize typical (inlier) observations, offering a structured approach to assessing epistemic uncertainty across different regions of the data distribution. We fit LOF on the first two t-SNE dimensions, assuming a contamination rate of $10\%$ (i.e., treating $10\%$ of the sample as outliers), and rank the top 150 outliers and inliers for analysis.
@@
Appendix B Predictive Distributions via Monte Carlo Dropout and Batch Normalization
Monte Carlo (MC) dropout was originally introduced as a regularization technique to prevent neural network overfitting [Srivastava et al., 2014]. However, it can also be used to approximate the predictive distribution by interpreting dropout as a variational Bayesian method. Indeed, Gal and Ghahramani [2016] demonstrated that applying dropout at both training and prediction corresponds to a variational approximation to the posterior over the network weights. Specifically, the dropout masks define a variational family where each mask corresponds to a Bernoulli-distributed perturbation of the network. From a practical standpoint, the predictive distribution is obtained by performing multiple stochastic forward passes with dropout enabled at prediction time [Izbicki, 2025]. Each pass applies a different dropout mask, resulting in varied outputs—in our case, different sets of means, variances, and weights of a Gaussian mixture for the same data point $\mathbf{x}$. A sample of $Y$ is then drawn from each configuration, producing a Monte Carlo sample from the predictive distribution $Y|\mathbf{x},D$.
Batch normalization (BN) [Ioffe and Szegedy, 2015] also introduces stochasticity, but through its reliance on mini-batch statistics during training. As shown by Teye et al. [2018], this stochasticity induces an implicit distribution over the network’s parameters, which can be interpreted as a form of approximate Bayesian prediction. During prediction, the batch statistics are fixed, and the model’s outputs can be treated as samples from an approximate posterior—similarly to the MC dropout case.
Appendix C Additional Results and Details for Real Data Experiments
@@
C.1 Evaluation Metrics
Let $\hat{R}(\cdot)$ denote a generic prediction interval. Given a test set $(\mathbf{X}_{1},Y_{1}),(\mathbf{X}_{2},Y_{2}),\ldots,(\mathbf{X}_{m},Y_{m})$, we evaluate performance using the following metrics:
@@ C.1 Evaluation Metrics
    $\text{AMC}=\frac{1}{m}\sum_{i=1}^{m}\mathbb{I}\left(Y_{i}\in\hat{R}(\mathbf{X}_{i})\right),$
@@
    where $\min\hat{R}(\mathbf{X})$ and $\max\hat{R}(\mathbf{X})$ represent the lower and upper bounds of the prediction interval, respectively, and $\alpha$ is the miscalibration level. The Interval Score Loss balances two key objectives: maintaining narrow prediction intervals while penalizing those that fail to cover $Y_{i}$, with larger penalties for greater deviations. By averaging these scores across all instances, AISL provides a measure that prioritizes the shortest interval while ensuring sufficient coverage. We chose the AISL following prior work in the literature, where it is commonly used as a summary metric [Rossellini et al., 2024], a paper that also addresses epistemic uncertainty in conformal predictions.
@@
    where $\mathbf{C}=(C_{1},\ldots,C_{m})$ represents a binary vector, with $C_{i}=\mathbb{I}(Y_{i}\in\hat{R}(\mathbf{X}_{i}))$ indicating whether the prediction interval $\hat{R}(\mathbf{X}_{i})$ covers $Y_{i}$, and $\mathbf{W}=(W_{1},\ldots,W_{m})$, where $W_{i}=\max\hat{R}(\mathbf{X}_{i})-\min\hat{R}(\mathbf{X}_{i})$. A strong correlation between coverage and interval width suggests a potential violation of conditional coverage, which requires their independence [Feldman et al., 2021]. However, $\rho=0$ does not guarantee conditional coverage, as non-adaptive methods like regression split can achieve zero correlation by maintaining constant-width intervals [Rossellini et al., 2024]. Thus, while this metric provides a useful proxy for assessing conditional coverage, it is not a definitive measure.
  • Outlier-to-inlier interval length ratio (ILR):
    $\text{ILR}=\frac{\frac{1}{|I_{\text{out}}|}\sum_{i\in I_{\text{out}}}\max\hat{R}(\mathbf{X}_{i})-\min\hat{R}(\mathbf{X}_{i})}{\frac{1}{|I_{\text{in}}|}\sum_{i\in I_{\text{in}}}\max\hat{R}(\mathbf{X}_{i})-\min\hat{R}(\mathbf{X}_{i})}\;,$
    where $I_{\text{out}}$ and $I_{\text{in}}$ denote the sets of outlier and inlier indices in the test set, respectively.Following the procedure described in Section A.2, we distinguish inliers from outliers by first reducing dimensionality with t-SNE [Van der Maaten and Hinton, 2008], and then applying the Local Outlier Factor (LOF) method [Breunig et al., 2000] using a contamination rate of $5\%$ (i.e., treating $5\%$ of the test instances as outliers). For this metric, we include all detected outliers in $I_{\text{out}}$ and define inliers $I_{\text{in}}$ as the 15% of instances with the lowest LOF scores. This metric captures how adaptively a method responds to epistemic uncertainty: higher ILR values indicate that outliers—typically found in data-sparse regions—are assigned wider prediction intervals relative to inliers, as expected from a well-calibrated uncertainty-aware method.
  • Average coverage on outliers (ACO):
    $\text{ACO}=\frac{1}{|I_{\text{out}}|}\sum_{i\in I_{\text{out}}}\mathbb{I}(Y_{i}\in\hat{R}(\mathbf{X}_{i}))\;,$
    where $I_{\text{out}}$ is the set of outlier indices in the test set, obtained using the same outlier detection procedure described above. This metric measures the proportion of outlier instances whose true response $Y_{i}$ falls within the predtion interval $\hat{R}(\cdot)$. Higher ACO values—ideally close to the nominal level $1-\alpha$ indicate that the method maintains reliable coverage even in data-sparse regions. As such, ACO complements the ILR metric by evaluating whether the adaptively wider intervals assigned to outliers are indeed well-calibrated.
C.2 Additional results
Table 5 summarizes the dataset details. Quantile regression results are reported in Tables 6, 7, 8, and 9, while standard regression results appear in Tables 10, 11, and 12. Additional metrics on outlier-to-inlier interval length ratios and outlier coverage are presented in Tables 9 (quantile regression), 13, and 14 (regression).
@@ C.2 Additional results
Table 9: Average coverage on outliers across methods and datasets in the quantile regression setting. This table reports the average prediction interval coverage for outlier observations—identified as data-sparse points—across 50 runs. Bold values indicate the best-performing method within a 95% confidence interval. EPICSCORE consistently delivers near-nominal coverage across most datasets.
Dataset EPIC-BART EPIC-GP EPIC-MDN CQR CQR-r UACQR-P UACQR-S airfoil 0.894 (0.029) 0.875 (0.032) 0.902 (0.024) 0.877 (0.031) 0.876 (0.032) 0.875 (0.033) 0.872 (0.031) bike 0.885 (0.009) 0.9 (0.011) 0.894 (0.011) 0.882 (0.01) 0.882 (0.01) 0.887 (0.009) 0.885 (0.009) concrete 0.824 (0.039) 0.855 (0.034) 0.875 (0.038) 0.845 (0.035) 0.844 (0.036) 0.878 (0.032) 0.845 (0.034) cycle 0.907 (0.009) 0.908 (0.01) 0.906 (0.011) 0.9 (0.009) 0.901 (0.009) 0.909 (0.01) 0.906 (0.009) electric 0.887 (0.01) 0.887 (0.009) 0.895 (0.009) 0.888 (0.009) 0.888 (0.009) 0.889 (0.01) 0.889 (0.009) homes 0.89 (0.007) 0.885 (0.007) 0.895 (0.008) 0.883 (0.007) 0.88 (0.007) 0.864 (0.009) 0.872 (0.007) meps19 0.904 (0.008) 0.903 (0.008) 0.899 (0.009) 0.897 (0.008) 0.897 (0.008) 0.9 (0.008) 0.897 (0.008) protein 0.893 (0.005) 0.895 (0.005) 0.896 (0.005) 0.894 (0.004) 0.894 (0.004) 0.894 (0.005) 0.894 (0.004) star 0.867 (0.016) 0.871 (0.017) 0.876 (0.016) 0.875 (0.018) 0.875 (0.018) 0.911 (0.021) 0.87 (0.018) superconductivity 0.863 (0.008) 0.872 (0.009) 0.879 (0.008) 0.913 (0.006) 0.912 (0.006) 0.932 (0.005) 0.912 (0.006) WEC 0.901 (0.006) 0.864 (0.008) 0.86 (0.009) 0.927 (0.005) 0.926 (0.005) 0.917 (0.005) 0.921 (0.005) winered 0.875 (0.029) 0.876 (0.03) 0.892 (0.026) 0.846 (0.032) 0.847 (0.032) 0.861 (0.033) 0.848 (0.032) winewhite 0.86 (0.015) 0.854 (0.016) 0.868 (0.014) 0.847 (0.016) 0.847 (0.016) 0.865 (0.02) 0.861 (0.016)
@@
Table 13: Average outlier-to-inlier interval length ratio across methods and datasets in the regression setting. The table reports the average ratio between prediction interval lengths for outliers and inliers, averaged over 50 runs. Bold values indicate the best-performing method within a 95% confidence interval. The interval length ration for Regression Split is omitted, as its constant interval length results will always return a ratio of 1. Overall, EPICSCORE consistently yields higher ratios across diverse datasets.
Dataset EPIC-BART EPIC-GP EPIC-MDN Mondrian Weighted airfoil 1.021 (0.039) 1.004 (0.014) 1.03 (0.035) 0.988 (0.023) 1.03 (0.045) bike 0.971 (0.014) 0.98 (0.006) 0.963 (0.02) 1.029 (0.018) 0.976 (0.016) concrete 1.05 (0.036) 1.03 (0.018) 1.11 (0.035) 0.99 (0.036) 1.137 (0.067) cycle 0.967 (0.007) 0.992 (0.002) 0.973 (0.008) 1.005 (0.006) 0.988 (0.005) electric 0.987 (0.008) 1.0 (0.003) 0.992 (0.01) 0.996 (0.005) 0.992 (0.008) homes 1.124 (0.019) 1.138 (0.015) 1.268 (0.035) 1.234 (0.026) 1.203 (0.035) meps19 1.047 (0.027) 1.057 (0.041) 1.097 (0.065) 1.099 (0.05) 1.074 (0.058) protein 0.999 (0.002) 0.996 (0.003) 1.008 (0.004) 1.003 (0.004) 1.0 (0.005) star 1.0 (0.007) 1.0 (0.001) 1.005 (0.007) 1.016 (0.011) 0.984 (0.025) superconductivity 1.129 (0.014) 1.061 (0.009) 1.107 (0.016) 1.211 (0.014) 1.111 (0.017) WEC 1.077 (0.007) 1.034 (0.005) 1.122 (0.015) 1.2 (0.014) 1.041 (0.007) winered 1.055 (0.019) 1.031 (0.009) 1.143 (0.027) 1.052 (0.02) 1.095 (0.045) winewhite 1.01 (0.007) 1.001 (0.001) 1.039 (0.012) 1.058 (0.013) 0.997 (0.011)
Table 14: Average coverage on outliers across methods and datasets in the regression setting. This table reports the average prediction interval coverage for outlier observations—identified as data-sparse points—across 50 runs. Coverage values closer to the nominal level of 0.9 reflect better adaptability to epistemic uncertainty. Bold values indicate the best-performing method within a 95% confidence interval. EPICSCORE consistently delivers near-nominal coverage across most datasets.
Dataset EPIC-BART EPIC-GP EPIC-MDN Mondrian Reg-split Weighted airfoil 0.896 (0.027) 0.883 (0.028) 0.894 (0.027) 0.91 (0.025) 0.902 (0.023) 0.903 (0.021) bike 0.895 (0.01) 0.901 (0.011) 0.897 (0.01) 0.909 (0.009) 0.903 (0.01) 0.895 (0.009) concrete 0.895 (0.029) 0.902 (0.027) 0.893 (0.028) 0.904 (0.023) 0.873 (0.033) 0.876 (0.029) cycle 0.89 (0.009) 0.901 (0.01) 0.892 (0.008) 0.914 (0.008) 0.901 (0.009) 0.898 (0.009) electric 0.9 (0.008) 0.898 (0.01) 0.902 (0.008) 0.909 (0.008) 0.906 (0.009) 0.911 (0.009) homes 0.869 (0.008) 0.873 (0.008) 0.896 (0.006) 0.894 (0.007) 0.847 (0.008) 0.878 (0.008) meps19 0.899 (0.007) 0.9 (0.007) 0.899 (0.008) 0.901 (0.012) 0.894 (0.008) 0.898 (0.007) protein 0.89 (0.005) 0.892 (0.005) 0.89 (0.005) 0.898 (0.004) 0.895 (0.004) 0.892 (0.004) star 0.892 (0.019) 0.899 (0.019) 0.906 (0.018) 0.899 (0.015) 0.885 (0.02) 0.879 (0.018) superconductivity 0.891 (0.007) 0.872 (0.008) 0.887 (0.008) 0.918 (0.005) 0.884 (0.007) 0.888 (0.006) WEC 0.861 (0.006) 0.85 (0.007) 0.871 (0.006) 0.897 (0.005) 0.85 (0.006) 0.852 (0.006) winered 0.87 (0.028) 0.841 (0.029) 0.881 (0.027) 0.858 (0.024) 0.843 (0.031) 0.846 (0.029) winewhite 0.872 (0.015) 0.876 (0.014) 0.878 (0.016) 0.9 (0.011) 0.863 (0.015) 0.867 (0.016)
C.3 Base model hyperparameters